# Certifying ground-state properties of quantum many-body systems

### v4

@@ B.2.5 Mirror symmetry
To summarize, for the 1D Heisenberg model, the SDP (18) with $d=4$ after reductions involves one positive semidefinite blocks of size $283r+$ 28, $N-1$ positive semidefinite blocks of size $273r+$ 27, and $3N$ positive semidefinite blocks of size $282r+$ 28; for the 2D Heisenberg model, the SDP (18) with $d=4$ after reductions involves one positive semidefinite blocks of size $1129L+$ 1, $L-1$ positive semidefinite blocks of size $129L$, and $3L$ positive semidefinite blocks of size $111L$.
@@ Appendix C Improving SDP bounds by imposing an extra positivity constraint
The bound given by the SDP relaxation for a fixed relaxation order can be improved by imposing an extra positivity constraint. We note that the constraint corresponding to the $k$-body physicality is
$\frac{1}{2^{k}}\left(1+\sum_{a_{1},\ldots,a_{k}}\langle\sigma_{1}^{a_{1}}\sigma_{2}^{a_{2}}\cdots\sigma_{k}^{a_{k}}\rangle\sigma_{1}^{a_{1}}\sigma_{2}^{a_{2}}\cdots\sigma_{k}^{a_{k}}\right)\geq 0,$ (32)
where $a_{i}\in\{0,x,y,z\}$, $i=1,\ldots,k$. As (32) is linear in the moments, we can add (32) to the constraints of (10). In theory, the larger $k$ is, the tighter the resulting bound is. However, a large $k$ also leads to a SDP of big size. In our experience, taking $k=8$ achieves a good balance between the increment of computational costs and the improvement of bounds.
@@ Appendix H Implementation details
All experiments were performed with one CPU core and 128G memory. In practice, we solve the dual problem of the SDP (18) (i.e., the sum-of-Hermitian-squares problem) since it contains less linear constraints and so can be solved more efficiently. We rely on the Julia package QMBCertify to create the optimization model and employ Mosek 10.0 as an underlying SDP solver. For the 1D model with $N=40,100$, it takes around 3.3h, 12h to generate a data point, respectively; for the 2D model model with $L=6,8,10$, it takes around 1.8h, 9.7h, 21h to generate a data point, respectively.

### v5

@@ B.2.5 Mirror symmetry
To summarize, for the 1D Heisenberg model, the SDP (18) with $d=4$ after reductions involves one positive semidefinite block of size $283r+$ 28, $N-1$ positive semidefinite blocks of size $273r+$ 27, and $3N$ positive semidefinite blocks of size $282r+$ 28; for the 2D Heisenberg model, the SDP (18) with $d=4$ after reductions involves one positive semidefinite block of size $1129L+$ 1, $L-1$ positive semidefinite blocks of size $129L$, and $3L$ positive semidefinite blocks of size $111L$.
@@ Appendix C Improving SDP bounds by imposing an extra positivity constraint
The bound given by the SDP relaxation for a fixed relaxation order can be improved by imposing an extra positivity constraint, namely that the $k$-body reduced density matrix $\rho_{[k]}$ is positive. This constraint can be written as:
$\rho_{[k]}=\frac{1}{2^{k}}\sum_{a_{1},\ldots,a_{k}}\langle\sigma_{1}^{a_{1}}\sigma_{2}^{a_{2}}\cdots\sigma_{k}^{a_{k}}\rangle\sigma_{1}^{a_{1}}\sigma_{2}^{a_{2}}\cdots\sigma_{k}^{a_{k}}\succeq 0,$ (32)
where $a_{i}\in\{0,x,y,z\}$, $i=1,\ldots,k$ (with $\sigma_{i}^{0}=\mathbf{1}$). As Eq.  (32) is linear in the moments, we can add it to the constraints of Eq.  (10). In general one expects that the larger $k$, the tighter is the resulting bound. However, as $\rho_{[k]}$ is a matrix of size $2^{k}\times 2^{k}$, a large $k$ also leads to a SDP of large size. In practice we notice that $k=8$ achieves a good balance between the computational cost and the improvement of bounds.
@@ Appendix H Implementation details
All experiments were performed with one CPU core and 128G memory. In practice, we solve the dual problem of the SDP (18) (i.e., the sum-of-Hermitian-squares problem) since it contains fewer linear constraints and so can be solved more efficiently. We rely on the Julia package QMBCertify to create the optimization model and employ Mosek 10.0 as an underlying SDP solver. For the 1D model with $N=40,100$, it takes around 3.3h, 12h to generate a data point, respectively; for the 2D model model with $L=6,8,10$, it takes around 1.8h, 9.7h, 21h to generate a data point, respectively.